you invest some money at 8% and half as much at 5%, you earn $3150 in annual interest, how much did you invest at each rate?
Answer by Flake(45)

Let X = money invested 8%
Let Y = monye invested at 5%
Given: Y = (1/2)X and total annual interest is $3150.
--Solution: Solve how much did you invest at each rate?
Interest = (money invested)*(interest rate)
So (8%)X + (5%)Y = $3150
0.08X + 0.05[(1/2)X] = $3150
0.105X = $3150
X = $30,000
Y = (1/2)X = $15,000